An open organ pipe has a length of 5 cm. (a) Find the fundamental frequency of vibration of this pipe. (b) What is the highest harmonic of such a tube that is in the audible range? Speed of sound in air is 340 m s-1 and the audible range is 20-20,000 Hz.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. Let the length of the resonating column will be = 1 Here V = 320 m/s Then the two successive resonance frequencies are (n + 1)/4 l and nv/4 l Here given (n + 1)v/4 l = 2592 ; λ = nv/4 l = 1944 ⇒ (n + 1)v/4 l - nv/4 l = 2592 – 19444 = 548 cm = 25 cm.